OpenStudy (anonymous):

in a video game,you earn 5 points for each gem you find. you earn 2 points each star you find. write and graph an equation that represents the number of gems and stars you must find to earn 250 points. What are 3 combinations of gems and stars you can find to earn 250 points.

OpenStudy (anonymous):

5g+2s = 250 and solve. Do you know how to solve?

OpenStudy (mathstudent55):

Let number of gems = g Let number of stars = s Let number of points = p p = 5g + 2s 5g + 2s = 250

OpenStudy (anonymous):

yeah do you get 48 gems and 5 stars?

OpenStudy (anonymous):

Yeah thats one way, there are multiple ways of solving, you want to isolate and solve for each as well.

OpenStudy (anonymous):

set either to zero and solve as well. so 5(0) + 2(x) = 250 and solve, and then 5(x) + 2(0) = 250

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

There are tons of variations of answers you can get. If you plot the points of the graph you would see.